NvidiaIntely (nvidiaintely):

Find the height of a rectangular prism if the surface area is 458.17 in2, the width is 6.2 in and the length is 8.3 in.

NvidiaIntely (nvidiaintely):

@Unavailabilityy

OpenStudy (jcreamer):

Using this formula (A = 2wl + 2lh + 2hw) to find surface area. 458.17=2*6.2*8.3+2h*8.3+2h*6.2 then solve for h

OpenStudy (jcreamer):

So 458.17=102.92+16.6h+12.4H or 458.17=102.92+29h

OpenStudy (jcreamer):

355.25=29h 12.25=h